QUESTION 5: (a) Innov Co. is a manufacturer of computer parts and sells parts in the local market, operating in the sub-sector of the information technology industry. Below is a summary of the financial information (dollars in millions) as at, and for the year ended 31 March 2018 for Innov Co. and five other companies the information technology industry, Ezy- Tech Ltd., Light Comm, Beta Parts, Decimal Inc., and T Zone Potential comparable companies Ezy-Tech Ltd Light Decima Innov Co. Beta Parts T Inc Comm Zone 2,508 2,004 Company value 592 1,810 3,512 Equity value Net operating assets (NOA) 2,366 603 1,657 3,250 1,812 405 112 499 352 757 454 Book value of equity Net operating profit after tax (NOPAT) 359 312 130 357 199 406 48 28 68 121 89 Net income (NPAT) Number of shares 31 57 48 95 78 outstanding (in millions) 120 70 250 150 350 300 Ezy-Tech Ltd., Beta Parts, Decimal Inc., and T Zone are manufacturers of computer parts. Only Decimal Inc. sells its products to foreign markets. Light Comm develops off the-shelf customer relationship management software for business clients and manufactures computer parts REQUIRED: (i) Identify the comparable companies for Innov Co. you will include in your valuation analysis using (1) the Price-to-NOPAT multiple, and (2) the Price-to- Book Value multiple. Provide justification for your selection. (3 marks) 40
(ii Based on your answers to part (i), compute the per share value of the equity for Innov Co. as at 31 March 2018 using (1) the Price-to-NOPAT multiple, and (2) the Price-to-Book Value multiple. (4 marks)
QUESTION 5 Continued: (b) Given the following information (dollars in millions) for Delta Limited and Omega Limited, determine the difference in the two companies' theoretically correct Price-to- Book Value ratio at the end of 2018. Delta Omega Limited Limited Information extracted from the financial statements for the year 2018 Net Operating Assets (NOA) Book value of equity Information estimated for the terminal period starting from 2019 $335 $285 $335 $285 RNOA (based on beginning balance of NOA) 15% 11% Expected ROE Weighted average cost of capital Expected growth rate 15% 11% 10% 10% 2% 2% (3 marks) (c) In market multiples analysis, there must be "consistency" between the multiple and the value driver. Explain briefly what is meant by this. (2 marks)
Answer #1

a-i). Based on the given information, Ezy-Tech, Beta Parts and T-Zone can be taken as comparable for Innov. Decimal Inc has a different market than Innov's and Light Comm does not manufacture computer parts so they are not suitable comparables.

Company Equity value (EV) Book value of equity (BV) NOPAT Price to NOPAT (EV/NOPAT) Price to BV (EV/BV)
Ezy-Tech                       603                       112                         28                   21.54                      5.38
Beta Parts                   1,657                       352                         62                   26.73                      4.71
T Zone                   1,812                       454                         89                   20.36                      3.99
Average                   22.87                      4.69

ii). A).Average price to NOPAT multiple = 22.87

Innov equity value = NOPAT*average multiple = 48*22.87 = 1,097.94 million

Shares O/S = 120 million

Price per share = Equity value/Shares O/S = 1,097.94/120 = $9.15 per share

B). Average price to book value mulitple = 4.69

Innov equity value = book value*average multiple = 312*4.69 = 1,464.58

Price per share = Equity value/Shares O/S = 1,464.58/120 = $12.20 per share

b). Using ROPI valuation model:

Formula Delta Omega
RNOA*Average NOA NOPAT                   50.25                   31.35
NOPAT - (NOA beg*WACC) ROPI                   16.75                      2.85
ROPI/(WACC -g) PV of ROPI                 209.38                   35.63
1 + (PV of ROPI/Owner's equity) PB ratio                      1.63                      1.13

Difference in PB ratio = 0.50

c). Consistency between the multiple and the value driver means that both the numerator and the denominator of the multiple are defined consistently. For example, in an EV to EBITDA multiple, both are measures of firm value while in a PB ratio, both are measures of equity. This consistency is important as it ensures that the multiples across companies can be compared as they are being computed using same value drivers.